Crew 188 Sol 3 Summary Report 31JAN2018
Summary Title: Winding down before ramping up
Author’s name: Ryan L. Kobrick, Ph.D., MDRS Crew 188 Commander Mission Status: All nominal
Sol Activity Summary:
Sol 2 merged into Sol 3 with the astronomical Super Blue Blood Moon event! The crew had varying levels of sleep hours leading into the early morning event. As you may have seen from earlier reports this was a very rare event, and the show met all expectations. All crew members were able to watch the Moon slowly disappear behind Earth’s shadow leaving a red glowing Moon. Looking at the red Moon created nice mental links to our Martian experience and submersion and let us ponder what it would be like to watch a dual-lunar event on Mars. Because of the rare event, today was dedicated to ramping up our science projects and resting for future EVAs. Personally I worked on my EVA metric study with the help of Tat to edit and name waypoints and record key data into a master spreadsheet. I also recorded a video podcast to post publicly (hopefully soon) interviewing Dr. Sarah Jane Pell. I am sure my students will find it very interesting and I’m expecting their follow up questions will be great! Renee and Zac started working on setting up the prototype of NASA Kennedy Space Center’s VEGGIE OASYS experiment in the Green Hab. It is great that the crew gets to work on an experiment that may one day be used on Mars, and super cool that it is already on the International Space Station. Thank you Dr. Luke Roberson for being the quarterback of this project at NASA KSC! Julia and Zac were also working in the Green Hab today, optimizing the layout and tending to the plants. Tat started a human factors analysis for the MDRS operational structure. This study could lead to consolidated task management and more efficient processes. The lighter day allowed the crew to sync files and prepare to share their experience. Several crewmembers sewed their mission patches onto their flight suits today, preparing to explore in style.
Ryan L. Kobrick, Ph.D.
MDRS Crew 188 Commander
Look Ahead Plan:
In the morning three crew members will head East to Phobos peak to examine human performance and capture 360 footage in 8k. The afternoon EVA will scout more roads and re-visit a previous site. The crew will be fairly busy with these two EVAs, but other projects will continue to unfold.
Anomalies in work:
MDRS Director is coordinating power system updates with contractors. There was a site visit today.
Robotic observatory currently not functional. MDRS Astronomy lead working problem.
The weather was clear all day and provided great views out the surrounding terrain.
Crew Physical Status:
Hydration and altitude adjustment seems to be going well. Some minor skin dryness for some crewmembers. Everyone seems more rested after a more relaxed day on campus.
Two EVAs are being planned and requested for tomorrow.
Reports to file:
1. EVA #4 and #5 requests
2. Ops Report
3. Sol Summary
4. Journalist Report
5. Creative Report
6. Green Hab Report
7. Astronomy Report
8. Daily Photos